# Recent Advances in Speech Language Models: A Survey

### v1

@@ 3.3.1 GAN-based Vocoder
For architectural choices, GAN-based vocoders focus on injecting inductive biases to generate audio waveforms. MelGAN Kumar et al. (2019) adds residual blocks with dilations in the generator to model the long-range correlation among the audio time steps and proposes a multi-scale architecture for the discriminator to model the different frequency ranges of the audio. Based on the idea of the multi-scale discriminator, HiFi-GAN Kong et al. (2020a) proposes a multi-period discriminator to model the diverse periodic patterns within the audio waveforms. To preserve high-frequency content, Fre-GAN Kim et al. (2021) employs the Discrete Wavelet Transform (DWT) to downsample and learn spectral distributions across multiple frequency bands. Unlike traditional approaches like Average Pooling (AP), DWT efficiently decomposes the signal into low-frequency and high-frequency sub-bands. BigVGAN Lee et al. (2022) introduces a periodic activation function called snake function along with an anti-aliased representation to reduce the high-frequency artifacts in the synthesized audio.
3.3.2 GAN-based Neural Audio Codec
Given that many neural audio codecs employ a GAN architecture, they can be effectively discussed within the context of GAN-based vocoders. Similar to its role as a tokenizer, although the primary objective of neural audio codecs is for audio compression, the encoded compact token sequences capture the essential information buried in the audio waveforms and therefore can be leveraged as a vocoder in SpeechLMs. EnCodec Défossez et al. (2022) uses a GAN architecture and proposes a novel generator including an encoder, a quantizer, and a decoder. The compressed audio representations are outputted by the quantizer by using Residual Vector Quantization (RVQ). Polyak et al. utilizes HiFi-GAN Kong et al. (2020a) as the vocoder backbone and proposes to disentangle the input features of a vocoder into distinct properties Polyak et al. (2021), which include semantic tokens, pitch tokens, and speaker embeddings. Such a design choice enables the codec to better perform on pitch and speaker-related tasks such as voice conversion and $F_{0}$ manipulation.
3.3.3 Other Types of Vocoder
@@
Pure Signal Processing Vocoder. Pure signal processing vocoders are traditional methods that rely on deterministic algorithms rather than deep learning models to synthesize speech Griffin and Lim (1984); Morise et al. (2016). However, this kind of vocoders introduces noticeable artifacts in the synthesized audio and is thus rarely used.
@@
Autoregressive Vocoder. Autoregressive vocoders generate audio waveforms one sample at a time, with each sample conditioned on the previously generated samples Oord et al. (2016). This approach allows for high-quality audio synthesis due to its sequential nature and the ability to capture intricate temporal dependencies within the audio signal. However, the sequential generation process can be computationally expensive and time-consuming, making autoregressive models less efficient compared to parallelized methods like GAN-based vocoders.
@@
Flow-based Vocoder. Flow-based vocoder aims to establish a series of invertible transformations that map a simple distribution, such as a Gaussian, to the complex distribution of audio samples. This mechanism allows for efficient sampling and density evaluation, enabling the model to synthesize audio in parallel rather than sequentially, which significantly enhances both speed and quality Prenger et al. (2019). Compared to GAN-based vocoders, Flow-based vocoders typically need more parameters and memory to train the model, which hinders them from being effectively utilized Kumar et al. (2019).
@@
VAE-based Vocoders. Variational Autoencoders (VAEs) are powerful generative models that learn to encode input data into a compressed latent space while allowing for the reconstruction of the original data Van Den Oord et al. (2017); Huang et al. (2019). However, VAE is seldom explored as the underlying architecture of vocoders.
@@
Diffusion-based Vocoder. Diffusion models have emerged in recent years as a powerful class of generative models that can be used for high-fidelity speech synthesis. They work by gradually adding noise to the input data (e.g. audio waveforms) to create a sequence of increasingly noisy representations, then learning to reverse this process to generate new samples Kong et al. (2020b); Chen et al. (2020); Lee et al. (2021). For instance, DiffWave Kong et al. (2020b) uses Denoising Diffusion Probabilistic Models (DDPM) to synthesize audio. Additionally, CosyVoice Du et al. (2024b) introduces a Conditional Flow-Matching (CFM) model that serves as a vocoder in TTS systems.
4 Training Recipes
Dataset Type Phase Hours Date
LibriSpeech Panayotov et al. (2015) ASR Pre-training 1k 2015
LibriLight Kahn et al. (2020) ASR Pre-training 60k 2019
People dataset Galvez et al. (2021) ASR Pre-training 30k 2021
VoxPopuli Wang et al. (2021) ASR Pre-training 1.6k 2021
Gigaspeech Chen et al. (2021) ASR Pre-training 40k 2021
Common voice Ardila et al. (2019) ASR Pre-training 2.5k 2019
WenetSpeech Zhang et al. (2022a) ASR Pre-training 22k 2022
LibriTTS Zen et al. (2019) TTS Pre-training 0.6k 2019
CoVoST2 Wang et al. (2020) S2TT Pre-training 2.8k 2020
CVSS Jia et al. (2022) S2ST Pre-training 1.9k 2022
Spotify podcasts Clifton et al. (2020) Podcast Pre-training 47k 2020
Fisher Cieri et al. (2004) Telephone conversation Pre-training 2k 2004
SpeechInstruct* Zhang et al. (2023a) Instruction-following Instruction-tuning - 2023
InstructS2S-200K* Fang et al. (2024) Instruction-following Instruction-tuning - 2024
@@ 4 Training Recipes
Table 2: A summary of popular datasets used in the pre-training and instruction-tuning phase of SpeechLMs. * means it is the speech version of the text dataset synthesized using TTS. S2ST and S2TT represent speech-to-speech translation and speech-to=text translation, respectively.
@@ 4.1 Features Modeled
The features modeled refer to the types of features outputted by the speech tokenizer and modeled by the language model component within a SpeechLM. These features play a crucial role in determining the capabilities and performance of SpeechLMs. Different features model the speech waveforms from different aspects. Based on recent developments, we can categorize the features modeled by SpeechLMs into two main types, including discrete features and continuous features.
4.1.1 Discrete Features
@@
Discrete features refer to quantized representations of speech signals that can be represented as distinct, countable units or tokens. These features are typically derived from speech signals through various encoding and quantization processes, resulting in a finite set of possible values. Discrete features are the most used features by SpeechLMs as they can be represented as tokens and be modeled exactly the same as the text tokens within a TextLM. The majority of speech tokenizers produce discrete tokens that better model the semantic information within a speech waveform (semantic understanding speech tokenizers, section 3.1.1), such as Chung et al. (2021); Hsu et al. (2021). This is because they primarily use understanding objectives such as MLM to model the contextual information of the waveforms when training the tokenizer. We refer to them as semantic tokens here.
@@
Most SpeechLMs only employ semantic tokens to represent speech. GSLM Lakhotia et al. (2021), the first-ever SpeechLM, compares three tokenizers, which include Contrastive Predictive Coding (CPC) Oord et al. (2018), wav2vec 2.0 Baevski et al. (2020), and HuBERT Hsu et al. (2021). It concludes that HuBERT performs the best on various tasks such as speech resynthesis and speech generation. A large number of works follow this setting and use HuBERT as the speech tokenizer Hassid et al. (2024); Nguyen et al. (2024); Zhang et al. (2023a). AudioPaLM Rubenstein et al. (2023) experiments the choice between w2v-bert Chung et al. (2021), USM-v1 Zhang et al. (2023c), and USM-v2 Rubenstein et al. (2023) (which is a modified version of USM-v1), and it concludes that USM-v2 is the best-performing speech tokenizer on ASR and Speech Translation (ST) tasks.
@@
Although semantic tokens excel at generating semantically meaningful speech because of the modeling of the contextual information within speech waveforms, researchers find out that the speech generated solely upon semantic tokens lacks expressive information such as prosody and different pitches or timbres Nguyen et al. (2023a, 2024). To conquer this limitation, paralinguistic tokens can be integrated into the modeling process to capture expressive information with speeches. Specifically, pGSLM Kharitonov et al. (2021) proposes to use the fundamental frequency (F0) and unit duration as prosody features to complement the HuBERT semantic tokens, and trains a multi-stream transformer language model to predict the semantic tokens, pitch (F0), and unit duration separately. Similarly, SPIRIT-LM Nguyen et al. (2024) complements the HuBERT semantic tokens with pitch and style tokens Duquenne et al. (2023). This incorporation of extra acoustic tokens allows SpeechLMs to more effectively capture expressive elements without significantly compromising semantic understanding Nguyen et al. (2024).
Another type is acoustic tokens, which are tokens aiming to capture the essential acoustic features to reconstruct high-fidelity speech, primarily obtained from neural audio codec models (see section 3.1.2). Codec models aim to learn the compressed representation of audio, so it is anticipated that both the semantic and acoustic information present in a speech waveform can be encoded in the representation. Some studies attempt to directly model the codec tokens in an autoregressive manner. VALL-E Wang et al. (2023b) utilizes codec tokens to achieve zero-shot TTS. It encodes a 3-second audio clip using EnCodec Défossez et al. (2022) as a prompt, enabling the TTS system to synthesize speech that matches the timbre information of the prompt. Viola Wang et al. (2024) uses codec tokens in a SpeechLM capable of performing ASR, TTS, and Machine Translation (in text).
@@
Discussion. Different types of tokens influence the speech quality of SpeechLMs in different ways, often resulting in trade-offs Borsos et al. (2023). For example, while semantic tokens align well with text and excel in producing semantically coherent speech, the generated speech often lacks acoustic details, such as high-frequency information. Recovering and enhancing these details typically requires postprocessing, like a diffusion model, which significantly increases the model’s latency. Conversely, acoustic tokens can facilitate the generation of high-fidelity audio but often struggle with inaccuracies in content generation Zhang et al. (2023b). Researchers have tried two ways to balance these trade-offs. The first involves combining semantic and acoustic tokens into a single sequence. AudioLM Borsos et al. (2023) proposes a hierarchical modeling scheme that first models semantic tokens from w2v-bert Chung et al. (2021) and then uses these tokens to predict acoustic tokens from SoundStream Zeghidour et al. (2021), which ultimately generates speech. However, this kind of approach increases sequence length, which increases modeling complexity. The second strategy leverages mixed tokens (see section 3.1.3) to jointly model semantic and acoustic information, showing promising results in Moshi Défossez et al. (2024) and SpeechGPT-Gen Zhang et al. (2024).
@@
Table 3: Four different methods of jointly modeling speech and text tokens.
4.1.2 Continuous Features
@@
Continuous features, in contrast to discrete features, are unquantized, real-valued representations of speech signals that exist on a continuous scale. These features capture fine-grained, nuanced aspects of speech that may be lost in discretization processes. Continuous features can include spectral representations like mel-spectrograms or latent representations extracted from neural networks. The exploration of leveraging continuous features to condition SpeechLMs is still in its infancy. Spectron Nachmani et al. (2023) performs speech continuation by predicting the spectrograms frame-by-frame. However, the generation of speech spectrograms still needs to be conditioned on text transcripts, which is not an end-to-end speech generation approach. Mini-Omni Xie and Wu (2024) extracts intermediate representations from a frozen Whisper encoder as input for the SpeechLM, whereas LauraGPT Du et al. (2023) employs an audio encoder trained alongside the SpeechLM to derive latent representations from input speech.
4.2 Training Stages
@@
Training a SpeechLM involves training the three main components: speech tokenizer, language model, and vocoder. Similar to TextLMs, the key to training SpeechLMs lies in effectively modeling speech continuation, which is primarily the responsibility of the language model. The speech tokenizer and vocoder usually rely on established methods and are trained using distinct training datasets specific to each SpeechLM approach. Therefore, This section reviews the main techniques used to train the language model component.
Following TextLMs, the training process for SpeechLMs can be divided into three stages: pre-training, instruction-tuning, and alignment. However, to our knowledge, there is currently no research specifically focused on the alignment process following instruction tuning. Therefore, we only discuss the works related to the pre-training and instruction-tuning stages of SpeechLMs.
4.2.1 Language Model Pre-training
@@
The pre-training of the language model in SpeechLMs is a critical phase that significantly influences the model’s ability to generate coherent and contextually relevant speech. This phase typically involves training the language model to autoregressively predict the next token on a large corpus of speech tokens. The primary objective during this stage is to learn the statistical patterns and dependencies inherent in the speech data, enabling the model to predict the next token in a sequence based on the preceding context. Table 2 includes popular datasets used in pre-training stage of SpeechLMs.
@@
Training data. SpeechLMs pre-training mainly leverages large-scale open-sourced speech data. Commonly used datasets include those for ASR Panayotov et al. (2015); Kahn et al. (2020); Galvez et al. (2021); Wang et al. (2021), TTS Zen et al. (2019), ST Jia et al. (2022); Wang et al. (2021), podcasts Clifton et al. (2020), and dialogues Cieri et al. (2004). Some datasets consist solely of speech data, while others include both speech and corresponding text transcripts. The inclusion of text transcripts can enhance the model’s representation by allowing it to learn the relationship between spoken language and its written form, which will be discussed later.
@@
Cold Initialization. Some SpeechLMs use cold initialization during the pre-training phase, where model parameters are initialized randomly. The pioneering SpeechLM—GSLM Lakhotia et al. (2021)—trained a transformer Vaswani (2017) from scratch to serve as the language model. This study demonstrated the effectiveness of the SpeechLM pipeline and compared performance across various speech tokenizer options. They found that HuBERT Hsu et al. (2021) outperformed CPC Oord et al. (2018) and wav2vec 2.0 Baevski et al. (2020) in understanding speech content and generating natural speech. SUTLM Chou et al. (2023) also uses a transformer as the language model. They studied the critical problem of jointly modeling speech and text tokens by comparing four different modeling methods: speech-only, text-only, concatenated speech-text, and alternating (interleaving) speech-text. They showed that the setting of alternating speech-text performs the best in cross-modal evaluations. Table 3 illustrates the four modeling methods.
@@
Some works leverage a different architecture from the standard transformer. Since there are no existing checkpoints for those self-proposed architectures, it is necessary to train them from scratch. For example, pGSLM Kharitonov et al. (2021) proposes a multi-stream transformer language model (MS-TLM) that takes multiple streams of input and predicts multiple streams of output to generate speech units, duration, and pitch embeddings simultaneously. dGSLM Nguyen et al. (2023b) introduced a dialogue transformer language model (DLM) to jointly model the dialogue speech data from the two speakers. To enable the listening ability of SpeechLMs while speaking, LSLM Ma et al. (2024) proposes to attach a streaming self-supervised learning (SSL) Encoder to an autoregressive token-based TTS Model. Viola Wang et al. (2024) introduced a multi-task auto-regressive codec language model to autoregressively generate codec tokens instead of speech unit tokens.
@@
Continued Pre-Training. In contrast to cold initialization, continued Pre-Training involves initializing the language model with pre-trained weights from a TextLM and then adapting it to handle speech tokens. This approach leverages the linguistic knowledge embedded in TextLMs, allowing for more efficient and effective training of SpeechLMs. Research by Hassid et al. (2024) found that starting with a textually pre-trained language model (OPT Zhang et al. (2022b) and LLaMA Touvron et al. (2023a)) can enhance the model’s convergence rate and significantly improve its speech understanding capabilities. They also demonstrated that while training from text-pretrained checkpoints outperforms cold initialization, training from image-pretrained checkpoints yields poorer results compared to cold initialization. This indicates that not all pre-trained checkpoints are equally effective. Additionally, AudioPaLM Rubenstein et al. (2023) trained the SpeechLM using PaLM and PaLM-2 Chowdhery et al. (2023); Anil et al. (2023b), showing that the SpeechLM benefits from both an increased size of the pre-trained checkpoint and a larger training dataset.
@@
The performance of SpeechLMs can be further enhanced by aligning the text and speech modality representations. SPIRIT-LM Nguyen et al. (2024) found that continually pretraining on TextLM checkpoints using interleaving text and speech tokens can significantly boost the model’s performance on speech understanding and generation. Additionally, their visualizations demonstrate that the similarity between text and speech features is notably higher in models trained with interleaved token sequences compared to those trained without this approach. AudioChatLlama Fathullah et al. (2024) aims to ensure that the model produces consistent outputs regardless of whether the input is text or speech. They address this challenge by treating text data in ASR datasets as prompts, allowing LLaMA to generate the corresponding responses. Consequently, both text and speech versions of the prompt can be utilized to train the model to provide the appropriate response. Spectron Nachmani et al. (2023) solves the text-speech representation alignment problem by jointly supervising multiple objectives. Specifically, the input speech prompt is first transcribed into its text tokens, and then the model predicts the text token response. Finally, the text response is synthesized to output speech.
4.2.2 Language Model Instruction-Tuning

### v2

@@ III-C1 GAN-based Vocoder
For architectural choices, GAN-based vocoders focus on injecting inductive biases to generate audio waveforms. MelGAN [89] adds residual blocks with dilations in the generator to model the long-range correlation among the audio time steps and proposes a multi-scale architecture for the discriminator to model the different frequency ranges of the audio. Based on the idea of the multi-scale discriminator, HiFi-GAN [34] proposes a multi-period discriminator to model the diverse periodic patterns within the audio waveforms. To preserve high-frequency content, Fre-GAN [91] employs the Discrete Wavelet Transform (DWT) to downsample and learn spectral distributions across multiple frequency bands. Unlike traditional approaches like Average Pooling (AP), DWT efficiently decomposes the signal into low-frequency and high-frequency sub-bands. BigVGAN [90] introduces a periodic activation function called snake function along with an anti-aliased representation to reduce the high-frequency artifacts in the synthesized audio.
III-C2 GAN-based Neural Audio Codec
Given that many neural audio codecs employ a GAN architecture, they can be effectively discussed within the context of GAN-based vocoders. In contrast to speech tokenizers, the decoder in the codec is leveraged as the vocoder [36, 22]. Polyak et al. [35] utilizes HiFi-GAN [34] as the vocoder backbone and proposes to disentangle the input features of a vocoder into distinct properties [35], which include semantic tokens, pitch tokens, and speaker embeddings. Such a design choice enables the codec to better perform on pitch and speaker-related tasks such as voice conversion and $F_{0}$ manipulation.
III-C3 Other Types of Vocoder
@@
Pure Signal Processing Vocoder. Pure signal processing vocoders are traditional methods relying on deterministic algorithms rather than deep learning models to synthesize speech [92, 93]. However, this kind of vocoder introduces noticeable artifacts in the synthesized audio and is rarely used.
@@
Autoregressive Vocoder. Autoregressive vocoders generate audio waveforms one sample at a time, with each sample conditioned on the previously generated samples [31]. This approach allows for high-quality audio synthesis due to its sequential nature and the ability to capture intricate temporal dependencies within the audio signal. However, the sequential generation process can be computationally expensive and time-consuming, making autoregressive models less efficient compared to parallelized methods like GAN-based vocoders.
@@
Flow-based Vocoder. Flow-based vocoder aims to establish a series of invertible transformations that map a simple distribution, such as a Gaussian, to the complex distribution of audio samples. This mechanism allows for efficient sampling and density evaluation, enabling the model to synthesize audio in parallel rather than sequentially, which significantly enhances both speed and quality [33]. Compared to GAN-based vocoders, Flow-based vocoders typically need more parameters and memory to train the model, which hinders them from being effectively utilized [89].
@@
VAE-based Vocoders. Variational Autoencoders (VAEs) are powerful generative models that learn to encode input data into a compressed latent space while allowing for the reconstruction of the original data [60, 94]. However, VAE is seldom explored as the underlying architecture of vocoders.
@@
Diffusion-based Vocoder. Diffusion models have emerged in recent years as a powerful class of generative models that can be used for high-fidelity speech synthesis. They work by gradually adding noise to the input data (e.g. audio waveforms) to create a sequence of increasingly noisy representations, then learning to reverse this process to generate new samples [95, 96, 97]. For instance, DiffWave [95] uses Denoising Diffusion Probabilistic Models (DDPM) to synthesize audio.
IV Training Recipes
@@
TABLE III: A summary of popular datasets used in the pre-training and instruction-tuning phase of SpeechLMs. * means it is the speech version of the text dataset synthesized using TTS. S2ST and S2TT represent speech-to-speech translation and speech-to=text translation, respectively.
Dataset Type Phase Hours Year
LibriSpeech [98] ASR Pre-Training 1k 2015
Multilingual LibriSpeech [99] ASR Pre-Training 50.5k 2020
LibriLight [100] ASR Pre-Training 60k 2019
People dataset [101] ASR Pre-Training 30k 2021
VoxPopuli [102] ASR Pre-Training 1.6k 2021
Gigaspeech [103] ASR Pre-Training 40k 2021
Common Voice [104] ASR Pre-Training 2.5k 2019
VCTK [105] ASR Pre-Training 0.3k 2017
WenetSpeech [106] ASR Pre-Training 22k 2022
LibriTTS [107] TTS Pre-Training 0.6k 2019
CoVoST2 [108] S2TT Pre-Training 2.8k 2020
CVSS [109] S2ST Pre-Training 1.9k 2022
VoxCeleb [110] Speaker Identification Pre-Training 0.4k 2017
VoxCeleb2 [111] Speaker Identification Pre-Training 2.4k 2018
Spotify Podcasts [112] Podcast Pre-Training 47k 2020
Fisher [113] Telephone conversation Pre-Training 2k 2004
SpeechInstruct* [8] Instruction-following Instruction-Tuning - 2023
InstructS2S-200K* [11] Instruction-following Instruction-Tuning - 2024
VoiceAssistant-400K* [10] Instruction-following Instruction-Tuning - 2024
@@ IV-A Features Modeled
The features modeled refer to the types of features or tokens outputted by the speech tokenizer and modeled by the language model component within a SpeechLM. These features play a crucial role in determining the capabilities and performance of SpeechLMs. Different features model the speech waveforms from different aspects. In this section, we summarize commonly used features in SpeechLMs and focus on how different features affect the performance of SpeechLMs. Based on recent developments, we can categorize the features modeled by SpeechLMs into two main types: discrete features and continuous features.
IV-A1 Discrete Features
@@
Discrete features (or discrete tokens) refer to quantized representations of speech signals that can be represented as distinct, countable units or tokens. These features are typically derived from speech signals through various encoding and quantization processes, resulting in a finite set of possible values. Discrete features are the most used features by SpeechLMs as they can be represented as tokens and be modeled exactly the same as the text tokens within a TextLM.
@@
Most SpeechLMs only employ semantic tokens (generated by semantic understanding tokenizers, Section III-A1) to represent speech, as semantic information plays the most crucial role in spoken communication. GSLM [37], the first-ever SpeechLM, compares three tokenizers, which include Contrastive Predictive Coding (CPC) [16], wav2vec 2.0 [17], and HuBERT [20]. It concludes that HuBERT performs the best on various tasks such as speech resynthesis and speech generation. A large number of works follow this setting and use HuBERT as the speech tokenizer [38, 5, 8]. AudioPaLM [39] experiments the choice between w2v-bert [19], USM-v1 [23], and USM-v2 [39] (a modified version of USM-v1), and it concludes that USM-v2 is the best-performing speech tokenizer on ASR and Speech Translation (ST) tasks.
@@
Although semantic tokens excel at generating semantically meaningful speech because of the modeling of the contextual information within speech waveforms, researchers find out that the speech generated solely upon semantic tokens lacks expressive information such as prosody and different pitches or timbres [114, 5]. To conquer this limitation, paralinguistic tokens can be integrated into the modeling process to capture expressive information with speeches. pGSLM [42] proposes to use the fundamental frequency (F0) and unit duration as prosody features to complement the HuBERT semantic tokens, and trains a multi-stream transformer language model to predict the semantic tokens, pitch (F0), and unit duration separately. Similarly, SPIRIT-LM [5] complements the HuBERT semantic tokens with pitch and style tokens [115]. This incorporation of extra acoustic tokens allows SpeechLMs to more effectively capture expressive elements without significantly compromising semantic understanding [5].
Another type is acoustic tokens, which are tokens aiming to capture essential acoustic features to reconstruct high-fidelity speech, primarily obtained from neural audio codec models (Section III-A2). Some studies directly model the codec tokens in a language model, which is often regarded as Codec Language Models (CodecLMs). For example, Viola [43] trains a CodecLM capable of performing ASR, TTS, and Machine Translation. Parrot [45] trains on VQ-VAE [60] tokens for modeling dual-channel spoken dialogue data.
@@
Discussion. Different types of tokens influence the speech quality of SpeechLMs in different ways, often resulting in trade-offs [81]. For example, while semantic tokens align well with text and excel in producing semantically coherent speech, the generated speech often lacks acoustic details, such as high-frequency information. Recovering and enhancing these details typically requires post-processing, like a diffusion model, which significantly increases the model’s latency. Conversely, acoustic tokens can facilitate the generation of high-fidelity audio but often struggle with inaccuracies in content generation [21]. Researchers have tried two ways to balance these trade-offs. The first involves combining semantic and acoustic tokens into a single sequence. AudioLM [81] proposes a hierarchical modeling scheme that first models semantic tokens from w2v-bert [19] and then uses these tokens to predict acoustic tokens from SoundStream [22], which ultimately generates speech. However, this kind of approach increases sequence length, which increases modeling complexity. The second strategy leverages mixed tokens (Section III-A3) to jointly model semantic and acoustic information, showing promising results in Moshi [9] and SpeechGPT-Gen [46].
TABLE IV: Four different methods of modeling speech and text tokens.
@@
IV-A2 Continuous Features
@@
Continuous features (or continuous tokens), in contrast to discrete features, are unquantized, real-valued representations of speech signals that exist on a continuous scale (continuous tokens). Continuous features can include spectral representations like mel-spectrograms or latent representations extracted from neural networks. Spectron [47] performs speech continuation by predicting the spectrograms frame-by-frame. Mini-Omni [10] and SLAM-Omni [41] extract intermediate representations from a frozen Whisper encoder as input for the SpeechLM, whereas LauraGPT [49] employs an audio encoder trained alongside the language model to derive latent representations from input speech. Continuous features can capture fine-grained, nuanced aspects of speech that may be lost in discretization processes. However, utilizing these features often necessitates modifying the off-the-shelf training pipeline of language models, as traditional text-based models are built to handle discrete units. Moreover, continuous features demand more storage capacity compared to their discrete counterparts.
IV-B Training Stages
@@
Training a SpeechLM involves training the three main components: speech tokenizer, language model, and vocoder. Similar to TextLMs, the key to training SpeechLMs lies in effectively modeling speech continuation, which is primarily the responsibility of the language model. The speech tokenizer and vocoder usually rely on established methods and are trained using distinct training datasets specific to each SpeechLM approach. Therefore, This section reviews the main techniques used to train the language model component. Following TextLMs, we divide the training process of SpeechLMs into three stages, including pre-training, instruction-tuning, and post-alignment.
IV-B1 Language Model Pre-Training
@@
The pre-training of the language model in SpeechLMs is a critical phase that significantly influences the model’s ability to generate coherent and contextually relevant speech. This phase typically involves training the language model to autoregressively predict the next token on a large corpus of speech tokens. The primary objective during this stage is to learn the statistical patterns and dependencies inherent in the speech data, enabling the model to predict the next token in a sequence based on the preceding context.
@@
Training data. SpeechLMs pre-training mainly leverages large-scale open-sourced speech data. Commonly used datasets include those for ASR [98, 100, 101, 102], TTS [107], ST [109, 102], podcasts [112], and dialogues [113]. Table III includes popular datasets used in the pre-training stage. Some datasets consist solely of speech data, while others include both speech and corresponding text transcripts. The inclusion of text transcripts can enhance the model’s representation by allowing it to learn the relationship between spoken language and its written form, which will be discussed later.
@@
Cold Initialization. Some SpeechLMs use cold initialization during the pre-training phase, where model parameters are initialized randomly. The pioneering SpeechLM—GSLM [37]—trained a transformer [25] from scratch to serve as the language model. This study demonstrated the effectiveness of the SpeechLM pipeline and compared performance across various speech tokenizer options. They found that HuBERT [20] outperformed CPC [16] and wav2vec 2.0 [17] in understanding speech content and generating natural speech. SUTLM [50] also uses a transformer as the language model. They studied the critical problem of jointly modeling speech and text tokens by comparing four different modeling methods: speech-only, text-only, concatenated speech-text, and alternating (interleaving) speech-text. They showed that the setting of alternating speech-text performs the best in cross-modal evaluations. Table IV illustrates the four modeling methods.
@@
Some works leverage a different architecture from the standard transformer. These models are typically trained from scratch when the architecture deviates from a standard transformer or TextLM too much. For example, pGSLM [42] proposes a multi-stream transformer language model (MS-TLM) that takes multiple streams of input and predicts multiple streams of output to generate speech units, duration, and pitch embeddings simultaneously. dGSLM [4] introduced a dialogue transformer language model (DLM) to jointly model the dialogue speech data from the two speakers. To enable the listening ability of SpeechLMs while speaking, LSLM [51] proposes to attach a streaming self-supervised learning (SSL) Encoder to an autoregressive token-based TTS Model.
@@
Continued Pre-Training. In contrast to cold initialization, continued Pre-Training involves initializing the language model with pre-trained weights from a TextLM and then adapting it to handle speech tokens. This approach leverages the linguistic knowledge embedded in TextLMs, allowing for more efficient and effective training of SpeechLMs. Hassid et al. [38] found that starting with a textually pre-trained language model (OPT [3] and LLaMA [26]) can enhance the model’s convergence rate and significantly improve its speech understanding capabilities. They also demonstrated that while training from text-pretrained checkpoints outperforms cold initialization, training from image-pretrained checkpoints yields poorer results compared to cold initialization. This indicates that not all pre-trained checkpoints are equally effective. Additionally, AudioPaLM [39] trained the SpeechLM using PaLM and PaLM-2 [116, 117], showing that the SpeechLM benefits from both an increased size of the pre-trained checkpoint and a larger training dataset.
@@
The performance of SpeechLMs can be further enhanced by aligning the text and speech modality representations. Some works align the text and speech representations in a single sequence. SPIRIT-LM [5] found that continually pretraining on TextLM checkpoints using interleaving text and speech tokens can significantly boost the model’s performance on speech understanding and generation. Additionally, their visualizations demonstrate that the similarity between text and speech features is notably higher in models trained with interleaved token sequences compared to those trained without this approach. Spectron [47] solve the text-speech representation alignment problem by jointly supervising multiple objectives. Specifically, the input speech prompt is first transcribed into its text tokens, and then the model predicts the text token response. Finally, the text response is synthesized to output speech. SpeechGPT [8] also adopts this concept but applies it during the instruction-tuning phase. Some other methods perform multi-sequence representation alignment. This approach simultaneously generates the text sequence and speech sequence(s). For example, Llama-Omni uses the LLM output hidden state to decode text tokens and generate discrete speech tokens simultaneously. Mini-Omni [10] generates a single sequence of text tokens and seven sequences of acoustic tokens in parallel, all aligned at the sentence level. Similarly, Moshi [9] generates one sequence of text tokens, one sequence of semantic tokens, and seven sequences of acoustic tokens in parallel, which are aligned at the word level.
Discussion. The primary goal of aligning text and speech representations is to leverage the strengths of text-based models to enhance speech-based models. Researchers have found that training a SpeechLM is significantly more challenging than training a TextLM. This difficulty arises because text serves as a concentrated form of knowledge, while speech requires models to independently learn the rules of spoken language. Aligning text and speech representations has demonstrated effectiveness, but it involves various trade-offs. First, text primarily conveys semantic information, which can improve a SpeechLM’s semantic modeling capabilities but may compromise its ability to capture paralinguistic features, such as tone and emotion, during alignment. Second, there are two main inference approaches for the aligned models: text-present and text-independent. Text-present inference decodes text and speech simultaneously, which may increase latency but enhances the SpeechLM’s reasoning abilities [10] and reduces possible hallucinations [9]. Conversely, text-independent inference is more efficient but may lack stability. Furthermore, the question of whether to incorporate text modality to enhance SpeechLM performance remains an open question, especially considering that humans typically acquire spoken language skills before mastering written language.
IV-B2 Language Model Instruction-Tuning